OpenStudy (anonymous):

Write an expression in simplest form for the volume of a box with a length of 5ab, a width of 3a, and a height of 2ab. A. 30a3b2 B. 7ab + 3a C. 10a3b3

OpenStudy (anonymous):

A box is a rectangular parallelepiped, it's volume is given by the product of length with width and height so here, the volume is \( 5ab \times 3a \times 3ab = 30 a^3b^2 \), so the correct option is \( \boxed{A} \)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Edit: \( 5ab \times 2a \times 3ab = 30 a^3b^2 \)